OpenStudy (anonymous):

r=12ft NEED ANSWER AND EXPLANATION

OpenStudy (annon):

what?

OpenStudy (anonymous):

r=12ft

OpenStudy (annon):

u mean pi*r^2?

OpenStudy (anonymous):

finding the area

OpenStudy (annon):

ya pi*(r^2)

OpenStudy (anonymous):

pi = 3.14 a = 3.14 * (12)^2 a = 3.14 * (12 * 12) a = 3.14 * 144 a = 452.12